# Minchul Kim, Anil Jain, Xiaoming Liu

### v2

@@ 6.5 Interpretability
Figure 18: Comparison between traditional face recognition (FR) pipelines and future paradigms integrating multimodal biometrics, reasoning, and explanations. Traditional FR systems output a simple match score based on feature similarity, whereas future systems will reason across modalities, dynamically assess uncertainty, and collaborate with humans through interpretable feedback loops.
Beyond static descriptions, emerging research explores chain-of-thought prompting [184] to enable step-by-step reasoning over input features, providing deeper multi-stage narratives. In face recognition, a system could first note matching high-level features (e.g., facial structure) before explaining divergences in fine details (e.g., minor pose variation). Incorporating such language-based explanations can significantly improve transparency and accountability in FR applications. Future work should focus on grounding these explanations tightly to visual evidence, minimizing hallucination, and ensuring faithfulness even as models scale to large, foundation model-based face recognition systems.
What becomes increasingly important is not just producing a match score—but doing so with calibrated confidence and explainability. We can begin to envision systems that, in uncertain cases, simulate multiple identity hypotheses, ask follow-up questions, request additional evidence, or provide probabilistic reasoning behind their decisions. Fig. 18 illustrates this conceptual shift, contrasting traditional FR pipelines with emerging interactive, reasoning-based biometric systems designed for calibrated confidence and human collaboration. This marks a shift from fully automated systems to a more collaborative model, where humans and deep learning systems coexist in a feedback loop. What is previously missing is the model’s ability to know when to intervene.
A critical aspect of enabling this loop is enhancing the model’s ability to know when to initiate interaction or request intervention. While existing research in Face Image Quality Assessment (FIQA) provides valuable groundwork by estimating input quality or recognizability [185, 186, 154], the next step involves integrating these signals into broader reasoning frameworks. Moving forward, understanding when and how to involve a human—with clear reasoning—will be a critical step. To enable this, we’ll also need new evaluation metrics—ones that go beyond accuracy, and reflect trust, interpretability, and decision quality in real-world scenarios.
@@ 7 Summary
Over the past half-century, automated face recognition has transformed remarkably, moving from early geometric and handcrafted methods to sophisticated deep learning models that often surpass human performance on benchmarks. This paper chronicled this journey, emphasizing the deep learning revolution driven by innovations in network architectures (ResNets to ViTs), specialized loss functions (like margin-based softmax), and massive datasets (e.g., WebFace42M), which have yielded current SoTA capabilities.
Despite technical successes, evidenced by near-saturation on high-quality benchmarks (LFW [26], CFP-FP [27]) and strong results on challenging ones (IJB-C [24]) leading to widespread deployments (from mobile face unlock and payment to airport security and immigration), face recognition is not solved. Significant hurdles remain for real-world deployments, including accuracy at massive scales (billions of identities), robustness to degraded imagery (low quality, pose, occlusion) from CCTV videos, model interpretability and data privacy for public trust. The limitations of face-only data drive interest in multimodal biometrics, while synthetic data generation offers potential solutions but faces challenges like domain gaps and identity capacity.
Despite remarkable progress and widespread adoption of face recognition (FR) technologies, several key challenges remain. Distinguishing between identical twins is a significant limitation, especially for national-scale systems, given the U.S. twin birth rate of 31.2 per 1,000 live births (CDC, 2022). While FR systems claim robustness over decades, longitudinal performance can degrade, requiring re-enrollment. Current models rely on embeddings rather than human-interpretable features, which limits explainability. Additionally, there are risks where facial data is used beyond its original intent, raising privacy and ethical concerns. Ensuring template integrity, encrypted matching, and resistance to spoofing attacks are persistent challenges. Balancing recognition accuracy and latency is critical, especially for billion-scale deployments. Addressing these issues is essential for advancing FR responsibly.
Looking ahead, research can focus on overcoming these persistent challenges. Key directions include developing robust multimodal systems for person recognition (not just multiple biomteric traits but also demographic and geolocation data), advancing explainable AI for transparency (e.g., evidence in courts of law for conviction), pushing synthetic data boundaries for realism and capacity, and evaluating the role of foundation models. Addressing large-scale gallery performance, calibrated confidence, and human-AI collaboration are also crucial next steps. As face recognition technology proliferates, continued innovation must be carefully coupled with responsible development and ethical considerations. These include use consent for data capture and data sharing, maintaining data security, and using data only for the purpose for which it was collected.

### v3

@@ 6.6 Interpretability
Figure 18: Comparison between traditional FR pipelines and future paradigms integrating multimodal biometrics, reasoning, and explanations. Traditional FR systems output a simple match score based on feature similarity, whereas future systems can reason across modalities, dynamically assess uncertainty, and collaborate with humans through feedback loops.
As interpretability advances, the next challenge is to pair match scores with calibrated confidence and clear reasoning. Future FR systems may, in uncertain cases, simulate multiple identity hypotheses, request additional evidence, or express probabilistic justifications. Fig. 18 illustrates this shift from static pipelines to interactive, reasoning-based systems that foster human collaboration. A key step toward such feedback loops is enabling models to determine when to involve humans. Building on Face Image Quality Assessment (FIQA) research that estimates input quality or recognizability [boutros2023cr, ou2024clib, chai2023recognizability], future work should integrate these signals into broader reasoning frameworks and develop metrics that go beyond accuracy to capture trust, interpretability, and decision quality.
6.7 Fairness and Bias in Face Recognition
As FR moved from research to practical deployment, ensuring equitable performance across demographic groups became critical. Fairness concerns intensified in the late 2010s following incidents of wrongful arrests [hill2020wrongful], biased airport screening [garvie2019america], and misidentification of public figures [snow2018amazon]. Since 2019, NIST’s FRVT includes demographic analyses [grother2019face]. For an extensive review, see Kotwal and Marcel [kotwal2025review]; this section summarizes key sources of bias, datasets, metrics, and mitigation strategies.
Sources of Disparities. Bias arises from multiple factors. Training data imbalance has long been cited [klare2012face, krishnapriya2019understanding], though disparities persist even in balanced sets [gwilliam2021rethinking]; Skin reflectance affects accuracy under poor lighting [cook2019demographic], while image quality and illumination improvements reduce group gaps [krishnapriya2020understanding]. Appearance factors such as hairstyle, facial hair, makeup, and occlusion drive gender differences more than gender itself [kurz2022exploring].
Datasets and Evaluation. Fairness evaluation requires demographically annotated datasets. Public options include RFW [wang2019racial], BFW [robinson2020face], BUPT and BalancedFace [wang2018racial].
Metrics and Mitigation. Fairness metrics captures shifts in similarity score distributions. NIST FRVT [grother2019face] contains large-scale demographic performance variations. Fairness Discrepancy Rate (FDR) quantifies these disparities, with values near one indicating greater parity [terhorst2020fairness]. Mitigation occurs through (1) preprocessing via demographic augmentation or synthetic data generation [kortylewski2019analyzing], (2) using adaptive architecture, losses or adversarial debiasing [wang2018racial, mitigating-face-recognition-bias-via-group-adaptive-classifier, gong2020debface], and (3) postprocessing with subgroup calibration or score normalization [robinson2020face].
6.8 Interconnected Areas with FR
Face Image Quality Assessment (FIQA). Recognition reliability depends strongly on image quality, influenced by pose, blur, and illumination. FIQA models estimate a recognizability score to guide quality aware training and inference. FIQA methods are either unsupervised, predicting quality from recognition feature certainty [Babnik2022FaceQAN, Babnik2023DifFIQA, meng2021magface, improving-face-recognition-with-a-quality-based-probabilistic-framework], or supervised, deriving pseudo quality labels through characterization and training an independent model [boutros2023cr, HernandezOrtega2019FaceQnet, Ou2021SDDFIQA, a-quality-guided-mixture-of-score-fusion-experts-framework-for-human-recognition]. FIQA now underpins FR by quantifying uncertainty and improving reliability with low quality inputs.
Presentation Attacks and Spoof Detection (PAD). As FR systems expand to consumer and border uses, they face attacks like printed photos, replayed videos, and three dimensional masks. PAD defends using texture cues, depth, or multi sensor fusion [liu2022spoof]. Progress in image forensics and deepfake detection adds augmentation [Li2020FaceXRay, Li2018WarpingArtifacts], frequency analysis [Liu2021SpatialPhase, Luo2021HighFreq], and disentanglement learning [Yan2023UCF, Yang2019HeadPose, hierarchical-fine-grained-image-forgery-detection-and-localization]. Interpretable detectors such as DDVQA BLIP [Zhang2024DDVQA] or M2F2Det [guo2025rethinking] employ vision language models for explanations, enhancing transparency.
Privacy Preserving Face Recognition (PFR). With FR in sensitive contexts, privacy protection is crucial. Cryptographic techniques use homomorphic encryption or secure multiparty computation for similarity computation without exposing raw faces [Erkin2009PETS, Yang2023Sensors], but computationally heavy. Transform based methods conceal by modifying representations. Early works use obfuscation [Kevenaar2005AutoID], while modern ones use adversarial or diffusion methods [Boutros2023ICCV, Huang2023CVPR].
6.9 Regulatory and Policy Perspectives.
Global scrutiny has turned fairness and accountability into regulatory imperatives. The GDPR [eu_gdpr_2016, zaborska2019biometric] classifies facial imagery and other biometric data as special personal data, requiring explicit consent or anonymization of major FR datasets. The EU AI Act [eu_ai_act_2024] designates biometric identification as high risk, mandating transparency, bias evaluation, and human oversight. Meanwhile, cities like San Francisco and Boston restrict public FR use, while others expand it for security purposes. These measures embed fairness as both scientific and regulatory.
@@ 7 Summary
Over the past fifty years, FR has advanced from geometric and handcrafted methods to deep learning models surpassing human accuracy. Progress in architectures (ResNets, ViTs), loss functions (margin based softmax), and large datasets (e.g., WebFace42M) has driven state of the art performance. Despite success on benchmarks (LFW [lfw], CFP FP [cfpfp], IJB-C [ijbc]) and deployment in authentication and security, challenges persist: scalability to billions, degraded imagery, limited interpretability, data privacy and fairness.
Future work should pursue multimodal and explainable systems, realistic synthetic data, and foundation models, emphasizing large scale robustness, confidence calibration, and ethical use through consent, security, and purpose limitation. Recent developments of FR datasets in the community have highlighted growing attention to the ethical and legal implications of datasets that lack clear and informed user consent, with increasing efforts to promote transparency and responsible data sourcing.
@@
[Uncaptioned image] Minchul Kim received his Ph.D. degree in Computer Science and Engineering from Michigan State University, East Lansing, MI, USA. He is currently a software engineer at Google, where he works on machine learning and computer vision applications. His research interests include face recognition, biometrics, and deep learning. During his Ph.D., he published in top-tier conferences and journals in the field of computer vision and biometrics.
[Uncaptioned image] Anil K. Jain () is a University Distinguished Professor in the Department of Computer Science and Engineering at Michigan State University. His research interests include pattern recognition, computer vision, and biometric authentication. Jain is a member of the U.S. National Academy of Engineering, the Indian National Academy of Engineering, the World Academy of Sciences, and the Chinese Academy of Sciences.
[Uncaptioned image] Xiaoming Liu () is a MSU Foundation Professor, and Anil and Nandita Jain Endowed Professor in the Department of Computer Science and Engineering at Michigan State University. He received his Ph.D. from Carnegie Mellon University in 2004. His research interests span computer vision, machine learning, and biometrics. He is an Associate Editor for IEEE Transactions on Pattern Analysis and Machine Intelligence. He is a fellow of IEEE and IAPR.